Unless otherwise stated, all objects are located near the Earth's surface, where g = 9.80 m/s2 .
In an emergency stop to avoid an accident, a shoulder-strap seatbelt holds a 62 kg passenger in place.

If the car was initially traveling at 86 km/h and came to a stop in 5.6 s along a straight, level road, what was the magnitude of the average force applied to the passenger by the seatbelt?
Answer #1

Force = rate of change of momentum
Force = m * Δv/t

Initial Velocity = 86km/h = 23.9 m/s
Final Velocity = 0

Δv = final velocity - initial velocity = 0 - 23.9 = - 23.9 m/s
-ve sign signifies velocity is falling.
rate of change of velocity = - 23.9 / 5.6 = 4.26 m/s^2

Force = m*a = 62* 4.26
Force = 264.12 N

Average force applied to the passenger by the seatbelt, F = 264.12 N